OpenStudy (anonymous):

Find the derivative of y=(cosX)^3

OpenStudy (anonymous):

is this \[y=\cos^3(x)\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if so \[y'=3\cos^2(x)\times -\sin(x)=-3\cos^2(x)\sin(x)\] by the "chain rule"

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

yw